the radius of a tepee is 15ft and the slanted sides of the tepee are 25 ft long, how high will the center of the tepee be.
Answer by edjones(8007)

a^2+b^2=c^2
a^2+15^2=25^2
a^2+225=625
a^2=400
a=20' height of the center.
.
Ed